3 attempts left Complete each nuclear equation. ( Give all nuclei in a. Rb 2o e 30
Answer #1

a)

Let the product be X

Balance mass number on both sides

mass of Rb = mass of e + mass of X

1*77 = 1*0 + M

M = 77

Balance atomic number on both sides

atomic of Rb = atomic of e + atomic of X

1*37 = 1*1 + A

A = 36

Atomic number 36 is for element Kr

So, the product is

77-Kr

Answer:

77

Kr

36

b)

Let the product be X

Balance mass number on both sides

mass of No = mass of He + mass of X

1*251 = 1*4 + M

M = 247

Balance atomic number on both sides

atomic of No = atomic of He + atomic of X

1*102 = 1*2 + A

A = 100

Atomic number 100 is for element Fm

So, the product is

247-Fm

Answer:

247

Fm

100

c)

Let the product be X

Balance mass number on both sides

mass of Cu = mass of e + mass of X

1*66 = 1*0 + M

M = 66

Balance atomic number on both sides

atomic of Cu = atomic of e + atomic of X

1*29 = 1*-1 + A

A = 30

Atomic number 30 is for element Zn

So, the product is

66-Zn

Answer:

66

Zn

30